OpenStudy (anonymous):

A circle has its center at (-2, 5) and a radius of 4 units. What is the equation of the circle?

jimthompson5910 (jim_thompson5910):

The general equation of a circle is (x-h)^2+(y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

where the center is (h,k) and the radius is r

jimthompson5910 (jim_thompson5910):

So what does that mean for us in this problem?

OpenStudy (anonymous):

(x-(-2))^2+(y-5)^2=4^2?

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

you can then rewrite x-(-2) as what?

OpenStudy (anonymous):

x+2+y-5=16 thats the whole equation simplified

jimthompson5910 (jim_thompson5910):

not sure what happened to your squares, but it should be (x+2)^2 + (y-5)^2 = 16

OpenStudy (anonymous):

oh, i forgot to add those in

jimthompson5910 (jim_thompson5910):

but it looks like you have the right idea

OpenStudy (anonymous):

thank you for the help

jimthompson5910 (jim_thompson5910):

you're welcome